# ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their ability to understand both natural language and code, driving their use in tasks like natural language-to-code (NL2Code) and code summarization. However, LLMs are prone to hallucination—outputs that stray from intended meanings. Detecting hallucinations in code summarization is especially difficult due to the complex interplay between programming and natural languages. We introduce a first-of-its-kind dataset with $\sim$10K samples, curated specifically for hallucination detection in code summarization. We further propose a novel Entity Tracing Framework (ETF) that a) utilizes static program analysis to identify code entities from the program and b) uses LLMs to map and verify these entities and their intents within generated code summaries. Our experimental analysis demonstrates the framework’s effectiveness, leading to a 73% F1 score. This approach provides an interpretable method for detecting hallucinations by grounding entities, allowing us to evaluate summary accuracy.
ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries
@@ 1 Introduction
Hallucination in natural language processing is defined as a condition in which a language model produces a text that is either incoherent or does not faithfully represent the provided source input Ji et al. (2023). Similarly, in the context of code summarization, hallucination can be defined as a condition in which the generated summary does not accurately capture the intent and implementation details of the given input code.
@@
[ht!] public RowBuilder int16(String name){ ColumnInt16 column = new ColumnInt16(_columns.size(), name, _offset); _offset += column.length(); _columns.add(column); return this;} Summary: ……. This method is used to add a new column of data type int16 (16-bit integer) to the existing data structure. It creates a new ColumnInt16 object with the given name and size (16 bits), updates the offset value to accommodate the new column………… Example 1 (LLama3-70B): Confused Data Type
@@
Figure 1: Proposed Methodology: This diagram illustrates our end-to-end Entity Tracing Framework (ETF), which takes source code and a corresponding summary as input. First, we use code parsers to extract entities from the source code and employ large language models (LLMs) to identify entities from the summary. Next, we apply string-based heuristics to match entities from the summary to the code. Following this, an LLM verifies the accuracy of each entity’s description by cross-referencing the source code with relevant sentences in the summary. This process enables the localization of hallucinated content in the summary, ultimately enhancing its interpretability.
@@
Hallucination can originate from a combination of factors, where one common reason could be the misinterpretation of code entities. This misunderstanding can impact the model’s ability to interpret the intended functionality of the code, resulting in an inaccurate portrayal of its purpose. For instance, consider the Example 1, where the intention of the int16 java method is to create a new 16-bit integer column (ColumnInt16) with a specified name, update the position for the next column, add it to the list of columns, and then return the RowBuilder object. However, the generated explanation introduces a non-existent int16 datatype and proceeds to discuss the rest of the logic as if it were valid. This could mislead a novice Java developer into believing that an int16 datatype exists in Java. Furthermore, several large language models (LLMs) like LLaMA and Granite failed to detect this hallucination. One reason could be that int16 is a valid datatype in other programming languages such as C, C++, C#, and Go, causing both humans and LLMs to confuse it with learning from those languages. Similarly, in the example shown in Figure 1, the java method getJobID() takes “jobName" as an argument and simply returns -1. The summary generated by the model provides a detailed explanation, including how the method getJobID() connects to the database and attempts to retrieve the jobID using the given jobName. Additionally, the summary mentions it stores the “jobStatus" in a variable. Clearly, the generated summary has no supporting entities in the code for db access and the model is relying on the method name to hallucinate a plausible summary of the method.
@@
In this work, we study different factors that can lead to hallucination and list down a taxonomy to map the common causes easily. Noting a lack of datasets to reliably research this topic. Therefore, we create a first-of-its-kind dataset for studying hallucination in code summarization with 411 summaries generated by seven different large language models, broken into 9933 entity-level samples. This dataset consists of code and a corresponding summary describing the code. The annotation consists of: a) NER b) Entity Description Verification (c) Overall Summary Quality (not focusing on code completeness or conciseness). We then introduce a framework that evaluates the correctness of the generated summary. For this, we verify if the entities discussed in the summary are present in the code and correctly described in the summary. The framework leverages code parsers like javalang [^1] https://github.com/c2nes/javalang to list the different entities in the code snippet and prompt-based approaches to detect entities in summary. We note that detecting entities in the generated summary is more difficult due to the high degree of polysemy Tabassum et al. (2020). For example, entities like "list", "while", "if", etc, can be a code entity or natural language entities. This necessitates our reliance on large language models with high reasoning capabilities for detecting entities on the summary side. We then map the detected entities from the summary to code by using string-matching heuristics. The sentences with unmapped entities can be considered as ungrounded (source of extrinsic hallucination). For each mapped entity, we then have a tuple <code, entity, intent-related sentence>, where the intent-related sentence can be considered as the sentence in summary mentioning the entity. The final step is to verify each tuple from the summary for intrinsic hallucination to assess the correctness of the code summary. Our experiments demonstrate the importance of localizing entities in the summary for effective hallucination detection. Our contributions are :
@@
  • A taxonomy covering diverse reasons that might lead to hallucination in the code summarization (Figure 2).
@@
  • A novel dataset [^2] We plan to open source the Data and Code for studying hallucination detection in code summarization, featuring 411 summaries from 7 LLMs and 10K entity-level samples (Table 1).
@@
  • A first-of-its-kind approach for entity-level hallucination detection in code summarization inspired by the insights from human behaviour during code reviews leading to a performance of 73% F1 score (Table 2).
@@ 2 Related Work
Hallucination in Natural Langauge: Rawte et al. (2024); Sahoo et al. (2024) review recent advances in hallucination detection in natural language, emphasizing its practical significance. Recently, prompt-based methods (Arora et al. (2022), Manakul et al. (2023), Agrawal et al. (2023), Dhuliawala et al. (2023)) are being used to detect hallucinations in the text produced by LLMs. Xiao and Carenini (2022) and Zhang et al. (2022) attempt to address entity-level verification in natural language inputs. Both of these works involve improving the correctness of natural language summaries and do not discuss anything in the context of code. We note that most of the hallucination detection frameworks Manakul et al. (2023); Arora et al. (2022); Dhuliawala et al. (2023); Valentin et al. (2024); Rebedea et al. (2023) in natural language do not enforce reference text for grounding. In our setup of code summarisation, the generated summary has to be evaluated with respect to a reference text (the code snippet). Therefore, neccessiating an approach which could compare the code summary to the code snippet. Maynez et al., 2020; Ji et al., 2023 discuss further fine-graining of hallucination in natural language as intrinsic and extrinsic hallucination. More specifically, Intrinsic hallucination occurs when the given text contradicts the reference, while Extrinsic hallucination happens when the text cannot be verified against the reference. We use a similar convention in our paper.
@@
Hallucination in Code Generation: The code generation space has captured significant attention due to its practical significance in software development. Jiang et al. (2024) discusses recent developments in code generation and suggests the importance of addressing hallucination for improving the reliability of LLMs. Liu et al. (2024) studies hallucination in code generation and proposes a categorization that encompasses five categories of hallucinations based on the conflicting objectives and varying degrees of deviation observed in code generation. Tian et al., 2024; Agarwal et al., 2024; Spracklen et al., 2024 advanced the field with datasets and frameworks addressing hallucination in code generation. These studies highlight that while LLM-generated code may be syntactically correct and semantically plausible, it often fails to execute as intended or meet requirements.
@@
Despite progress in hallucination detection, code summarization remains underexplored. Kang et al., 2024 and Zhang, 2024 focused on inconsistencies in comment generation, addressing specific aspects like design constraints and parameter types, but their methods face challenges due to reliance on execution environments. In contrast, our approach validates the full functionality of generated outputs, independent of external dependencies, offering a more reliable solution by grounding entities in the input code and verifying their intent.
3 Datasets
@@
To create the hallucination dataset for code summarization, we consider code snippets from Java programming language and CodeXGLUE Lu et al. (2021) – Code-To-Text dataset. We focused on Java programming language due to its widespread relevance in the industry. It offers a rich set of entities (such as classes, methods, and variables) due to its structured design and strict typing system. The dataset was annotated by 8 annotators who are experts in Java and held at least a Master’s degree in Computer Science, with some having a PhD in the field. On average, the annotators had 4+ years of experience in Java programming. We report the statistics in Table 1 and describe the data curation process below:
@@
Hallucination Labeling: For each detected code entity in a summary, all sentences describing that entity are considered relevant. To account for the scenario where the relevant sentence can be noisy, we introduced a third label, "IRRELEVANT", which can be used to evaluate the performance of the intent-detection module and removed during the preprocessing. Thus, we obtain tuples of (code, entity, relevant sentences) for each entity. A total of 9933 such tuples, sampled from 441 summaries, were selected for human annotation (hired based on volunteering) to detect hallucinations. Out of these, 4354 tuples (from 222 summaries) were independently reviewed by two different sets of annotators, leading to a Cohen Kappa score of 0.72, implying high agreement. The conflicts were resolved by two independent meta-annotators. The annotators were asked to evaluate overall summary by assigning a label of ‘GOOD", ‘FAIR" and ‘POOR" We observe that, on average, 1.33 entities were marked as hallucinated for the summaries rated as ’FAIR" or ’POOR". Therefore, we consider a summary as hallucinated if at least one of the entities is hallucinated. After pre-processing, we consider the instance with labels "CORRECT" and "INCORRECT" in human data and treat the "IRRELEVANT" label predicted by the model as "INCORRECT". We provide the complete annotation guideline in Appendix E.2.
4 Categorization of Factors for Hallucination in Code Summarization
@@
Figure 2: Taxonomy of hallucination in code summarization based on the causes of hallucination. We start with four broad categories and then present a fine-grained classification of each category.
@@
In this section, we describe the various factors that could lead to hallucination in code summaries (Figure 2) based on what we learned from the annotation process. This classification, based on the underlying factors of hallucination, offers insights into the generative behaviors of language and code models. We describe these categories of hallucination factors below and discuss their statistical analysis in Figure 3.
@@
HC1: Based on Identifier Name Bias: Name Bias refers to the tendency of language models to rely on identifier names when interpreting code. We classify this bias into three subcategories based on its source: 1)variables, 2)functions, 3)libraries. The model can misinterpret code due to the linguistic characteristics of these entity names. As the semantics of the code is defined by the underlying logic rather than their lexical meaning of entities, this may lead to hallucination. In the example shown in Figure 1, the model (Granite-20B) incorrectly assumes that getJobID is about retrieving a job ID, based purely on their names, even though the actual code logic suggests otherwise.
@@
HC2: Insufficient knowledge: This involves scenarios where the model generates incorrect summaries due to lack of knowledge. This may include an incorrect explanation of the imported libraries that the model did not see in its training data, incorrect information about the keyword, etc. We further divide this category into two parts:
@@
2) Non-contextual code involves the scenario when the input does not contain the complete code and mentions an unseen library or an unknown construct whose functionalities are not understood by the model. For example, in the code sample shown in the HC2 Example, the model incorrectly describes the purpose of SQLException.
@@
HC3: Code Complexity: This pertains to the model’s tendency to produce incorrect code summaries due to high code complexity. This may stem from the model’s insufficient reasoning capabilities to understand the code or the misinterpretation of user instructions. Key factors contributing to complexity include: 1) Length: Longer code is more complex, requiring deeper understanding and involving more interdependencies with more potential points of failure. 2) Lexical Complexity: Complex vocabulary, including diverse operands and operations, increases the number of elements to track and understand. 3) Logical Complexity: Code with high cyclomatic complexity, multiple paths, or distant method invocations increases the challenge of comprehension.
[ht!] public <V>V execute(RedisCallback<V> cb) { Jedis jedis = jedisPool.getResource(); boolean success = true; try {return cb.execute(jedis);} catch (JedisException e) { success = false; if (jedis != null){ jedisPool.returnBrokenResource(jedis);} throw e; } finally { if (success) { jedisPool.returnResource(jedis);}}} Summary: …it then returns the ‘jedis" object to the pool using the ‘returnBrokenResource" method if the ‘success" variable is ‘false"… HC3: Codellama-7B confusion with conditions
In the code snippet shown in HC3 Example, the model Codellama-7b Roziere et al. (2023) produces an incorrect interpretation of the condition. This may be due to increased complexity due to nesting leading to more complicated logic that are challenging to understand by model.
HC4: Natural language Context: This refers to cases where natural language in code snippets, such as outdated comments or log statements, causes hallucinations in code summaries. In the code snippet shown in HC4 Example, the LLama3-70B model incorrectly infers that the property variable contains a list of key-value pairs inferred from a commented line. However, the ‘property" variable contains an alphanumeric string followed by one or more semi-colons.

### v4

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their ability to understand both natural language and code, driving their use in tasks like natural language-to-code (NL2Code) and code summarisation. However, LLMs are prone to hallucination—outputs that stray from intended meanings. Detecting hallucinations in code summarisation is especially difficult due to the complex interplay between programming and natural languages. We introduce a first-of-its-kind dataset, CodeSumEval, with $\sim$10K samples, curated specifically for hallucination detection in code summarisation. We further propose a novel Entity Tracing Framework (ETF) that a) utilises static program analysis to identify code entities from the program and b) uses LLMs to map and verify these entities and their intents within generated code summaries. Our experimental analysis demonstrates the framework’s effectiveness, leading to a 73% F1 score. The proposed approach provides a method for detecting hallucinations by tracing entities from the summary to the code, allowing us to evaluate summary accuracy and localise the error within the summary.
ETF: An Entity Tracing Framework for Hallucination Detection
in Code Summaries
@@ 1 Introduction
Hallucination in natural language processing is defined as a condition in which a language model produces a text that is either incoherent or does not faithfully represent the provided source input Ji et al. (2023). Similarly, in the context of code summarisation, hallucination can be defined as a condition in which the generated summary does not accurately capture the intent and implementation details of the given input code.
@@
[ht!] public RowBuilder int16(String name) { ColumnInt16 column = new ColumnInt16(_columns.size(), name, _offset); _offset += column.length(); _columns.add(column); return this; } Summary: ……. This method is used to add a new column of data type int16 (16-bit integer) to the existing data structure. It creates a new ColumnInt16 object with the given name and size (16 bits)……….. Example 1 (LLama3-70B): Confused Data Type
@@
Figure 1: Proposed Methodology: This diagram illustrates our end-to-end Entity Tracing Framework (ETF), which takes source code and a corresponding summary as input and returns whether the summary is hallucinated or not. First, we use code parsers to extract entities from the source code and employ large language models (LLMs) to identify entities from the summary. Next, we apply string-based heuristics to match entities from the summary to the code. Following this, an LLM verifies the accuracy of each entity’s description by cross-referencing the source code with relevant sentences in the summary. This process enables the localisation of hallucinated content in the summary, ultimately enhancing its explainability.
@@
Hallucination in code summaries often stems from misinterpreting entities, complex logic, or limited model understanding. This can impact the ability of models to interpret the intended functionality of the code, resulting in an inaccurate portrayal of its purpose. With the recent advances in language models Achiam et al. (2023); Touvron et al. (2023); Mishra et al. (2024); Team et al. (2023), the hallucination detection in code summarisation can become very challenging due to the tendency of models to produce convincing but inaccurate summaries. For instance, consider the Example 1, where the intention of the int16 java method is to create a new 16-bit integer column (ColumnInt16) with a specified name, update the position for the next column, add it to the list of columns, and then return the RowBuilder object. However, the generated explanation introduces a non-existent int16 datatype and proceeds to discuss the rest of the logic as if it were valid. It can be noted that int16 is a valid datatype in other programming languages such as C, C++, C#, and Go, causing LLMs to confuse it with learning from those languages. In general, these statements could also mislead a novice Java developer into believing that an int16 datatype exists in Java. Furthermore, several large language models (LLMs) like Llama and Granite failed to detect this hallucination. Similarly, in the example shown in Figure 1, the Java method getJobID() takes “jobName" as an argument and simply returns -1. The summary generated by the model provides a detailed explanation, including how the method getJobID() connects to the database and attempts to retrieve the jobID using the given jobName. Additionally, the summary mentions it stores the “jobStatus" in a variable. Clearly, the generated summary has no supporting entities in the code for database access, and the model is relying on the method name to hallucinate a plausible summary of the method.
We focus on detecting hallucinations in code summaries, cases where generated text misrepresents the code. Our study is motivated by the growing real-world use of LLM-based code documentation tools like Amazon Q Developer [^1] Amazon Q Developer and IBM Watsonx Code Assistant [^2] IBM Watsonx Code Assistant. Therefore, detecting hallucinations in code-related outputs is critical to ensuring the reliability and correctness of these systems in general software engineering workflows.
We study different factors that can lead to hallucination and list down a taxonomy to map the common causes easily. Noting a lack of datasets to reliably research this topic, we create a first-of-its-kind dataset, CodeSumEval, for studying hallucination in code summarisation with 411 summaries generated by seven different large language models, broken into 9933 entity-level samples. This dataset consists of code and a corresponding summary describing the code. The code summaries are generated with the original CodeXGLUE Lu et al. (2021) dataset, which consists of code snippets scraped from open-source GitHub repositories. All code snippets used are unaltered and authentic, thereby reflecting realistic, real-world usage scenarios. The annotation consists of a) NER, b) Entity Description Verification, and (c) Overall Summary Quality (not focusing on completeness or conciseness). Importantly, all the hallucination cases discussed in the paper are naturally occurring in code summaries from the models and are not synthetically constructed or perturbed, thereby incorporating the real-world challenges of convincing summaries.
@@
Further, we introduce a framework that evaluates the correctness of the generated summary. For this, we verify if the entities discussed in the summary are present in the code and correctly described in the summary. The framework leverages code parsers like javalang333javalang Repository to list the different entities in the code snippet and prompt-based approaches to detect entities in the summary. We note that detecting entities in the generated summary is more difficult due to the high degree of polysemy Tabassum et al. (2020). For example, entities like "list", "while", "if", etc, can be code entities or natural language entities. This necessitates our reliance on large language models with high reasoning capabilities for detecting entities on the summary side. We then map the detected entities from the summary to code by using string-matching heuristics. The sentences with unmapped entities can be considered as unground (source of extrinsic hallucination). For each mapped entity, we have a tuple <code, entity, intent-related sentence>, where the intent-related sentence can be considered as the sentence in summary mentioning the entity. Finally, we verify each tuple from the summary for intrinsic hallucination to assess the correctness of the code summary. Our experiments demonstrate the importance of localising entities in the summary for effective hallucination detection. Our contributions are:
@@
  • A taxonomy covering diverse reasons that might lead to hallucination in the code summarisation (Figure 2).
@@
  • A novel dataset CodeSumEval [^4] GitHub Link for studying hallucination detection in code summarisation, featuring 411 summaries from 7 LLMs and $\sim$ 10K entity-level samples (Table 1) with an explanation of causes for hallucinations as per taxonomy.
@@
  • A first-of-its-kind approach for hallucination detection in code summarisation inspired by the insights from human behaviour during code reviews, leading to a performance of 73% F1 score (Table 2).
@@ 2 Related Work
Hallucination in Natural Language: Rawte et al. (2024); Sahoo et al. (2024) review recent advances in hallucination detection in natural language, emphasising its practical significance. Recently, prompt-based methods (Arora et al. (2022), Manakul et al. (2023), Agrawal et al. (2023), Dhuliawala et al. (2023)) are being used to detect hallucinations in the text produced by LLMs. Xiao and Carenini (2022) and Zhang et al. (2022) attempt to address entity-level verification in natural language inputs. Both of these works involve improving the correctness of natural language summaries and do not discuss anything in the context of code. We note that most of the hallucination detection frameworks Manakul et al. (2023); Arora et al. (2022); Dhuliawala et al. (2023); Valentin et al. (2024); Rebedea et al. (2023) in natural language do not enforce reference text for grounding. In our setup of code summarisation, the generated summary has to be evaluated with respect to a reference text (the code snippet). Therefore, necessitating an approach which could compare the code summary to the code snippet. Maynez et al., 2020; Ji et al., 2023 discuss further fine-graining of hallucination in natural language as intrinsic and extrinsic hallucination. More specifically, Intrinsic hallucination occurs when the given text contradicts the reference, while Extrinsic hallucination happens when the text cannot be verified against the reference. We use a similar convention in our paper.
@@
Hallucination in Code Generation: The code generation space has captured significant attention due to its practical significance in software development. Jiang et al. (2024b) discusses recent developments in code generation and suggests the importance of addressing hallucination for improving the reliability of LLMs. Liu et al. (2024) studies hallucination in code generation and proposes a categorisation that encompasses five categories of hallucinations based on the conflicting objectives and varying degrees of deviation observed in code generation. Tian et al., 2024; Agarwal et al., 2024; Spracklen et al., 2024 advanced the field with datasets and frameworks addressing hallucination in code generation. These studies highlight that while LLM-generated code may be syntactically correct and semantically plausible, it often fails to execute as intended or meet requirements.
@@
Despite progress in hallucination detection, code summarisation remains underexplored. Kang et al., 2024 and Zhang, 2024 focused on inconsistencies in comment generation, addressing specific aspects like design constraints and parameter types, but their methods face challenges due to reliance on execution environments. In contrast, our approach validates the full functionality of generated outputs, independent of external dependencies, offering a more reliable solution by grounding entities in the input code and verifying their intent.
3 CodeSumEval Dataset
@@
To create the CodeSumEval dataset, we consider code snippets from the Java programming language and CodeXGLUE Lu et al. (2021) – Code-To-Text dataset. We focused on the Java programming language due to its widespread relevance in the industry. It offers a rich set of entities (such as classes, methods, and variables) due to its structured design and strict typing system. The dataset was annotated by 8 annotators who are experts in Java and hold at least a Master’s degree in Computer Science, with some having a PhD in the field. On average, the annotators had 4+ years of experience in Java programming. We report the statistics in Table 1 and describe the data curation process below:
@@
Hallucination Labelling: For each detected code entity in a summary, all sentences describing that entity are considered relevant. To account for the scenario where the relevant sentence can be noisy, we introduced a third label, "IRRELEVANT". These instances can be used to evaluate the performance of the intent-detection module and removed during the preprocessing. Thus, we obtain tuples of (code, entity, relevant sentences) for each entity. A total of 9933 such tuples, sampled from 441 summaries, were selected for human annotation (hired based on volunteering) to detect hallucinations. Out of these, 4354 tuples (from 222 summaries) were independently reviewed by two different sets of annotators, leading to a Cohen Kappa score of 0.72, implying high agreement. The conflicts were resolved by two independent meta-annotators. The annotators were asked to evaluate the overall summary by assigning a label of ‘GOOD’, ‘FAIR’, and ‘POOR’. We observe that, on average, 1.33 entities were marked as hallucinated for the summaries rated as ‘FAIR’ or ‘POOR’ . Therefore, we consider a summary as hallucinated if at least one of the entities is hallucinated. After preprocessing, we consider the instance with labels "CORRECT" and "INCORRECT" in human data and treat the "IRRELEVANT" label predicted by the model as "INCORRECT". We provide the complete annotation guideline in Appendix E.2.
4 Categorisation of Factors for Hallucination in Code Summarisation
@@
Figure 2: Taxonomy of hallucination in code summarisation based on the causes of hallucination. We start with four broad categories and then present a fine-grained classification of each category.
@@
In this section, we describe the various factors that could lead to hallucination in code summaries (Figure 2) based on what we learned from the annotation process. This classification, based on the underlying factors of hallucination, offers insights into the generative behaviours of language and code models. We describe these categories of hallucination factors below and discuss their statistical analysis in Figure 3.
@@
HC1: Based on Identifier Name Bias: Name Bias refers to the tendency of language models to rely on identifier names when interpreting code. We classify this bias into three subcategories based on its source: 1)Variables, 2)Functions, and 3)Libraries. The model can misinterpret code due to the linguistic characteristics of these entity names. As the semantics of the code is defined by the underlying logic rather than their lexical meaning of entities, this may lead to hallucination. In the example shown in Figure 1, the model (Granite-20B) incorrectly assumes that getJobID is about retrieving a job ID, based purely on their names, even though the actual code logic suggests otherwise.
@@
HC2: Insufficient knowledge: This involves scenarios where the model generates incorrect summaries due to the lack of knowledge. This may include an incorrect explanation of the imported libraries that the model did not see in its training data, incorrect information about the keyword, etc. We further divide this category into two parts:
@@
2) Non-contextual code: This involves the scenario when the input does not contain the complete code and mentions an unseen library or an unknown construct whose functionalities are not understood by the model. For example, in the code sample shown in the HC2 Example, the model incorrectly describes the purpose of SQLException.
@@
HC3: Code Complexity: This pertains to the model’s tendency to produce incorrect code summaries due to high code complexity. This may stem from the model’s insufficient reasoning capabilities to understand the code or the misinterpretation of user instructions. Key factors contributing to complexity include: 1) Length: Longer code is more complex and involves more interdependencies with more potential points of failure. 2) Lexical Complexity: Complex vocabulary, including diverse operands and operations, increases the number of elements to track and understand. 3) Logical Complexity: Code with high cyclomatic complexity, multiple paths, or distant method invocations increases the challenge of comprehension.
[ht!] public <V> V execute(RedisCallback<V> cb) { Jedis jedis=jedisPool.getResource(); boolean success = true; try { return cb.execute(jedis); } catch (JedisException e) { success = false; if (jedis != null) { jedisPool.returnBrokenResource(jedis); } throw e; } finally { if (success) { jedisPool.returnResource(jedis); } } } Summary: …it then returns the ‘jedis" object to the pool using the ‘returnBrokenResource" method if the ‘success" variable is ‘false" HC3: Codellama-7B confusion with conditions
In the code snippet shown in HC3 Example, the model Codellama-7b Roziere et al. (2023) produces an incorrect interpretation of the condition. This may be due to increased complexity due to nesting, leading to more complicated logic that is challenging for the model to understand.
@@
HC4: Natural language Context: This refers to cases where natural language in code snippets, such as outdated comments or log statements, causes hallucinations in code summaries. In the code snippet shown in HC4 Example, the LLama3-70B model incorrectly infers that the property variable contains a list of key-value pairs inferred from a commented line. However, the ‘property" variable contains an alphanumeric string followed by one or more semicolons.